OpenStudy (anonymous):

The variables x and y satisfy the equation x^{n}y = C, where n and C are constants. When x = 1.10,y = 5.20, and when x = 3.20, y = 1.05. (i) Find the values of n and C.

OpenStudy (anonymous):

\[x^{n}y = C\]

OpenStudy (experimentx):

two equation two variables. that should be solvable.

OpenStudy (anonymous):

I got to the point \[{1.10^n \over 3.20^n}=0.2019\]

OpenStudy (experimentx):

Thats even excellent. I was gong to say, take log on both sides, n ln x + ln y = ln c put values, you will have two linear equations on n and ln c, solve them and find values.

OpenStudy (anonymous):

So how do you get an answer from there?

OpenStudy (experimentx):

(1.10/3.20)^n = 2.019

OpenStudy (anonymous):

And from there?

OpenStudy (experimentx):

take log on both sides.

OpenStudy (experimentx):

n = ln(2.019)/ln(1.10/3.20)

OpenStudy (anonymous):

n=-0.658

OpenStudy (anonymous):

?

OpenStudy (experimentx):

could be, now find the value of C using this n.

OpenStudy (anonymous):

4.88=c?

OpenStudy (experimentx):

does it fit both data?? it it fits, then it's right ... else it's wrong.

OpenStudy (anonymous):

OK.

OpenStudy (anonymous):

I got 0.488 and 4.88??? weird..

OpenStudy (experimentx):

for which value

OpenStudy (anonymous):

for C

OpenStudy (experimentx):

sorry, i mislead you (1.10/3.20)^n = 2.019 should have been => (1.10/3.20)^n = 0.2019

OpenStudy (anonymous):

Yes, it works now. Thanks :)